OpenStudy (anonymous):

Solve this system algebraically. 9x + 2y = 5 y - 2x + 3 = 0 {(11/13, -17/13)} {(-11/13, -17/13)} {(-11/13, 17/13)} {(11/13, 17/13)}

OpenStudy (anonymous):

HELP PLEASE!!

OpenStudy (anonymous):

9x + 2y =0,5 Y - 2x + 3 = 0. 9x=-2y,x=-2/9y.5y-2((-2/9)y)+3=0,5y+4/9y=-3,49/9y=-3,y=-27/49.x=98/243.

mathslover (mathslover):

ok so first of all given to us are two equations : \[\large{1) \implies 9x+2y=5}\] \[\large{2) \implies y-2x+3=0}\] Solve for x in 1st equation : \[\large{9x+2y=5}\] \[\large{9x=5-2y}\] \[\large{x=\frac{5-2y}{9}}\]

OpenStudy (anonymous):

or 9x + 2y = 5 y - 2x + 3 = 0

mathslover (mathslover):

Now put this value of x in the second equation like : \[\large{y-2x+3=0}\] \[\large{y-2(\frac{5-2y}{9})+3=0}\] \[\large{y-\frac{(-10+4y)}{9}+3=0}\]

OpenStudy (anonymous):

it sounds complicated..

mathslover (mathslover):

@omfgitsshelby Can you tell me what you get after solving the above equation for y... ?

OpenStudy (anonymous):

its 0

OpenStudy (anonymous):

do you need an explination or you just need the answer

OpenStudy (anonymous):

hold on <3

OpenStudy (pradipgr817):

9x+2(2x-3)=5 13x=11 x=11/13, y=2x(11/13)-3=(22-39)/13=-17/13 x=11/13, y=-17/13

mathslover (mathslover):

take your time @omfgitsshelby

OpenStudy (anonymous):

@mathslover how exactly did you solve for x can you walk me threw that because i didn't understand how you got that

OpenStudy (anonymous):

rfl

OpenStudy (anonymous):

? @Algebraic!

OpenStudy (pradipgr817):

@omfgitsshelby i've solved it see

OpenStudy (anonymous):

i did thank you c:

OpenStudy (anonymous):

i saw is what i meant c: